OpenStudy (anonymous):

MEDAL AND FAN! Please help! :) In a bag of small balls 1/4 are green, 1/8 are blue, 1/12 are yellow and the remaining 26 white. How many balls are blue?

OpenStudy (whpalmer4):

If 1/4 are green, 1/8 are blue, and 1/12 are yellow, with the remaining 26 being white, that means that the white balls make up 1-(1/4+1/8+1/12) of the total number of balls.

OpenStudy (whpalmer4):

Or in mathematical terms, if \(x\) represents the number of balls, \[x*(1-(\frac{1}{4}+\frac{1}8+\frac{1}{12})) = 26\]Instead of doing the tedious work on the fractions, I would just multiply each term by the least common multiple of 4,8,12 which is 24. Remember, however, that \(x\) is the total number of balls, and you'll still have to figure how how many are blue.